OpenStudy (anonymous):

Please help! Ive been stucked in this question for quite a while. Heres the question:

OpenStudy (anonymous):

Given a^2+b^2+c^2=ab+bc+ca Prove a=b=c

RhondaSommer (rhondasommer):

we know that the following is true: (a-b)^2 + (b-c)^2 + (a-c)^2 >= 0 because each of the added terms is greater than or equal to zero. Multiplying out you get: a^2 - 2ab + b^2 + b^2 - 2bc + c^2 + a^2 - 2ac + c^2 >= 0 Collecting like terms and rearranging you get: 2a^2 + 2b^2 + 2c^2 - 2ab - 2bc - 2ac >= 0 Then dividing each side of the equation by 2 and rearranging you get: a^2 + b^2 + c^2 >= ab + ac + bc

RhondaSommer (rhondasommer):

does that help?

RhondaSommer (rhondasommer):

okeeeeey....bye i guess? give medal please :)

RhondaSommer (rhondasommer):

@mathmale does that make sense? i feel like it made a little but not a lot of sense...

RhondaSommer (rhondasommer):

we know that the following is true: \[ (a-b)^2 + (b-c)^2 + (a-c)^2 >= 0 \]because each of the added terms is greater than or equal to zero. Multiplying out you get: \[ a^2 - 2ab + b^2 + b^2 - 2bc + c^2 + a^2 - 2ac + c^2 >= 0 \] Collecting like terms and rearranging you get: \[2a^2 + 2b^2 + 2c^2 - 2ab - 2bc - 2ac >= 0 \] Then dividing each side of the equation by 2 and rearranging you get: \[a^2 + b^2 + c^2 >= ab + ac + bc\]

OpenStudy (anonymous):

Yayyy okay thanks Umm how do i give out a medal?

OpenStudy (anonymous):

But I need to prove a=b=c So wouldnt i need to prove a-b=0?

OpenStudy (anonymous):

Oh and wouldn't i know a^2+b^2+c^2>= 0 since x^2>=0 must be true if x is a real number? I wonder if i can get anything out of this: (a+b+c)^2=a^2+b^2+c^2+2(ab+bc+ca) Let a^2+b^2+c^2=x this: (a+b+c)^2=3x

OpenStudy (anonymous):

Ohhh ive looked it up. I know how to give out medals now lol. Thanks anyway

RhondaSommer (rhondasommer):

wow i got five medals :) sorry i was asleep so i didnt see you replies...do you still need help?